OpenStudy (anonymous):

Find the product: 2y^2(y^4-y^3-4y^2+4y+5)

OpenStudy (shadowfiend):

Just multiply it out. So: \[2y^2(y^4 - y^3 - 4y^2 + 4y + 5)\] You have to distribute the \(2y^2\) throughout the parenthesized part: \[2y^2 \cdot y^4 - 2y^2 \cdot y^3 - 2y^2 \cdot 4y^2 + 2y^2 \cdot 4y + 2y^2 \cdot 5\] Can you go ahead and do those multiplications?

OpenStudy (anonymous):

do i add or multiply the exponents?

OpenStudy (shadowfiend):

When you multiply two different powers of the same number (or variable), you add the exponents. You would multiply them if you double raised. So: \[3x^2 \cdot x^5 = 3x^{2 + 5} = 3x^7\] \[(x^2)^5 = x^{2\cdot 5} = x^{10}\]

OpenStudy (anonymous):

2y^2(y^4-y^3-4y^2+4y+5) 2y^6-2y^5-8y^6+8y^310y^2??

OpenStudy (shadowfiend):

I think you messed up your third term. The rest look good.